5x^2-2x-2=0 solve by completing the square
Answer by stanbon(75887)

5x^2-2x-2=0 solve by completing the square
--------
5(x^2-(2/5)+ ? = 2 + 5*?
----
5(x^2 -(2/5)x + (1/5)^2) = 2 + 5(1/5)^2
-------
5(x-(1/5))^2 = 11/5
-----
(x-(1/5)^2 = 11/25
-----
x-(1/5) = sqrt(11)/5 or x - (1/5)= -sqrt(11)/5
===============
x = (1+sqrt(11)/5 or x = (1-sqrt(11))/5
====================
Cheers,
Stan H.
====================